OpenStudy (anonymous):

What is the difference of negative 1/10 - negative 8/10 in the lowest terms?

OpenStudy (p0sitr0n):

-1/10 - (-8/10) = -1/10 + 8/10 = 7/10. This is irreducible since 7 and 10 are coprime.